# Epistemic diversity across language models mitigates knowledge collapse

### v1

@@ 3. Approach
We study the progress of model performance in AI ecosystems containing 1, 2, 4, and 16 models, each re-trained on their collective output over 10 iterations. All models begin as identical copies of a pre-trained model but are fine-tuned on unique subsets of the training data. This segmentation approach enables us to isolate and examine the effect of diversity across models on their performance under fixed model architecture and a fixed per-ecosystem resource budget for data size (number of tokens, $N$) and compute.
@@
In the first iteration step, the training data is identical to the generation dataset for a given model. Thereafter, it evolves in content because it is sourced from the outputs of the previous iteration: At each iteration, the $M$ generated datasets are concatenated, shuffled, and uniformly redistributed to the $M$ models for the next round. Throughout the ten iterations, the training and generation datasets for each model within an ecosystem are equal in size, $n=N/M$. For example, for two models ($D=M=2$), the sizes of the training and generation datasets are each half of the original size, $n=N/2$ (see Figure 2). In the case of the single-model ecosystem ($M=1$), our approach parallels the methods used in prior work on single-model collapse (e.g., Shumailov et al., 2024).
Translated to the real world, the segmentation approach broadly mirrors domain- or community-specific models (e.g., Feng et al., 2024b; Tanksley et al., 2025) trained on distinct segments of internet data. While tailored to different communities that might use different inputs, the generated outputs are collectively stored in a single data pool, from which future training samples are drawn.
@@ 3.1. Ecosystem diversity $D$ and model performance
Drawing from ecology, we use the HSD to measure ecosystem diversity, that is, diversity across models. The HSD quantifies the effective number of equally common elements (models, species, or categories) in a population (Roswell et al., 2021; Jost, 2006; Hill, 1973). Let $M$ denote the number of models, then we calculate HSD as:
@@
Because of equal weight across models ($w_{m}=1/M$), the diversity simplifies to the number of models in a given ecosystem:
@@
Consistent with prior work (Shumailov et al., 2024), we compute each model’s perplexity on the original Wikitext-2 test split to assess model performance. Perplexity reflects a model’s ability to complete the test set text sequences (lower is better) and indicates the degree of alignment between the model’s probability distribution $p_{\theta}$ and the true distribution $p_{data}$. The width of the perplexity distribution serves as an indicator of model collections’ expressivity, which we use in Figure 5.
@@
(3) $\mu_{t}=\frac{1}{M}\sum_{m=1}^{M}\mu_{m,t}(N_{test})\text{ , }$
whereas $\mu_{m,t}(N_{test})$ corresponds to the mean perplexity across the text sequences of the original Wikitext2 test split.
To compare ecosystems across iterations, we calculate an aggregated mean perplexity. Let $T$ denote the number of iteration steps, $\mu_{t}$ an ecosystem’s mean perplexity for a given iteration step $t$, then, the aggregated perplexity $\mu_{T}$ equals:
@@ 3.2. Generation and training details
We use GPT2 (Radford et al., 2019) and OPT-125m (Zhang et al., 2022), two open-accessible, small, and well-studied language models used in prior work on collapse (e.g., Shumailov et al., 2024; Herel and Mikolov, 2024).
Analogous to prior work (Herel and Mikolov, 2024; Shumailov et al., 2024), we use Wikitext2 (Merity et al., 2016) as the original dataset. To reduce computational effort during training and generation, we use a randomly sampled subset of two-fifths of the original size (=$N$), while keeping the original Wikitext2 validation and test splits for evaluation. Using a block size of 128 tokens, the sizes of the training–generation, validation, and test splits correspond to 7,500, 1,900, and 2,200 data points, respectively.
@@
All experiments are carried out using the HuggingFace Transformers library (Wolf et al., 2019) and the Python Lightning framework (Falcon and The PyTorch Lightning team, 2019).
@@ 4. Results
We conducted various experiments, testing the impact of increasing diversity on ecosystems of AI models trained on their collective output. Under a fixed data size and compute budget, we find that performance improves for increased diversity up to a certain point. In other words, we find an optimal degree of diversity greater than zero (see Figure 3).
Figure 4 shows the mean perplexity for the OPT-125m and GPT2 collections of 1, 2, 4, and 16 models trained on unique segments of the artificial training data across 10 re-training iterations. Ecosystems with a diversity of four ($D=M=4$) achieve the lowest aggregated mean perplexity for both model types, with the effect more pronounced for OPT-125m. Ecosystems with lower diversity ($M=1,2$) perform well in early stages but show rapid perplexity increases in later iterations.
The least diverse ecosystem ($D=1$) is the best performer only at $t=0$, when all models are trained on real data [^4] Progress, including perplexity for the original models, is shown in Figure 7 in the appendix.. The only difference is the per-model sample size ($n=N/M$). Conversely, the ecosystem with the highest diversity ($M=16$) begins with relatively high perplexity, but, after evolving, it approaches similarly low perplexity values as $M=4$. In addition to aggregated mean perplexity, diversity also affects the rate of perplexity change after 10 iterations. While low-diversity ecosystems exhibit a clearly positive rate, indicating a strong performance decay, the highest-diversity ecosystem yields a negative rate (see Figure 4).
Figure 5 shows the mixed perplexity distributions of the GPT2 and OPT-125m ecosystems on the original training dataset, indicating that collections with greater diversity exhibit both greater variety in the information represented and lower alignment with the true distribution.
Figure 3. AI ecosystems exhibit an optimal degree of diversity (Hill–Shannon Diversity $D$) for performance, operationalized as the mean perplexity on the original test set averaged across 10 training–generation iterations. For both model types, GPT2 and OPT-125m, this perplexity minimum occurs at $D=M=4$. The colored symbols correspond to the colors in Figure 2, which indicate different levels of diversity (e.g., orange corresponds to $D=2$).
@@
Figure 4. Diversity across models $D$ mitigate model collapse. Comparison of ecosystems’ perplexity with increasing diversity ($D=M$) at fixed dataset size across ten iteration steps. For both model types, $M=4$ yields the lowest mean perplexity at the early and late stages of evolution. After 10 iterations, the perplexity increasing rate is positive for low diversity ($M=1,2$), near 0 for $m=4$, and negative for the highest diversity ecosystem ($M=16$). The shaded region is a combined standard deviation across the models and data points.
(a) OPT-125m
(b) GPT2
Figure 5. Perplexity distributions widen with increasing diversity. Mixed perplexity distribution on the original Wikitext2 train data points, evaluated by all models in a given ecosystem with diversity ($D$ = $M$ = 1, 2, 4, 16) at iteration steps $t$ = 0 (left) and $t$ = 9 (right). As ecosystems’ diversity increases, the distributions become wider indicating greater expressivity. Simultaneously, their means shift toward higher perplexity, suggesting greater divergence from the distribution in the original training set. Across all ecosystems, the distributions narrow from the beginning to the end of the model iterations.
@@ 5. Theoretical interpretation
The inverted U-shaped relationship between diversity and performance reflects a trade-off between opposing effects of diversity across models. To understand these effects, we draw on the three sources of error in model collapse: statistical approximation, functional expressivity, and functional approximation (Shumailov et al., 2024).
The first error arises because the empirical distribution $\hat{p}_{data}$ of the training data does not accurately match the true distribution $p_{data}$ [^5] The probability distributions are illustrated in Figure 6 in the appendix.. In the first iteration step, when models are trained on subsets of the original, non-model-generated training dataset, this limited representativeness is solely due to finite sampling, and the error would vanish as the sample size approaches infinity. However, in subsequent iteration steps, the training data is no longer merely a subset of the original dataset but is increasingly contaminated with examples outside the support of the true distribution, generated in previous iterations. Consequently, the representativeness of a training set depends on both recall (the fraction of the true distribution covered by the training data) and precision (the fraction of training examples that actually come from the true distribution).
The second source of error concerns the models’ capacity to collectively represent the training data distribution $\hat{p}_{data}$. For illustration, consider fitting a mixture of two Gaussians with a single Gaussian in the case of a single model, compared to fitting the same mixture with two Gaussians in the case of a collection of two diverse models (example borrowed from Shumailov et al., 2024).
The third source of error concerns the models’ capacity to learn what they are capable of representing. In our experiments, this error can be neglected because all models share the same architecture and training setup.
When all models are trained on subsets of the original training dataset, we operate under fixed precision; therefore, recall decreases as $n$ shrinks with the number of models $M$. The resulting reduction in statistical approximation for models trained on smaller datasets explains the negative effect of diversity, particularly evident in the first iteration.
At the same time, because individual models are forced to specialize in subpopulations, increasing the number of models improves functional expressivity of their ensemble. The widths of perplexity distribution reflect the collections’ expressivity, illustrated in Figure 5. Improved expressivity can increase the representativeness (recall and precision) of the collectively generated data, thereby reducing the statistical approximation error even when individual models are trained on fewer data points. Conversely, once the quality of the artificially generated training data drops because its contamination with examples outside the support of the true distribution, increasing the sample size $n$ per model can become harmful for subsequent training (compare to Sorscher et al., 2023; Xie et al., 2023; Feng et al., 2024a). This dynamic likely explains the strong performance decay of the $M=1$ ecosystem toward higher number of iterations shown in Figure 4. Thus, the improved expressivity and its downstream benefits explain the positive effect of diverse collections on model performance.
In essence, less data reduces the statistical approximation capacity of diverse collections when training data quality is sufficiently high. Simultaneously, however, diversity increases expressivity, which can improve data quality in later iterations and thereby improve statistical approximation in subsequent steps.

### v2

@@ 3. Approach
We study the progress of model performance in AI ecosystems containing 1, 2, 4, and 16 models, each re-trained on their collective output over 10 iterations. All models begin as identical copies of a pre-trained model but are fine-tuned on equal-sized, non-overlapping subsets of the training data [^4] This approach using non-overlapping train sets is sometimes called bootstrap aggregation (bagging) (Dietterich, 2000; Ganaie et al., 2022) and has been applied in early work studying performance of model ecosystems (e.g., Krogh and Vedelsby, 1994). This segmentation approach enables us to isolate and examine the effect of diversity across models on their performance under fixed model architecture and a fixed per-ecosystem resource budget for data size (number of tokens, $N$) and compute. We first introduce the primary experimental setting, which demonstrates a clear positive effect of ecosystem diversity on model performance in a controlled, though limited, setting. Then, in Section 3.3, we present variations of this primary setting that aim to approximate more realistic conditions and to examine how the location of the optimum shifts with factors such as model size, data size, incorporation of real data into training, and temperature sampling. We denote the following main variables: $D$ (ecosystem diversity), $M$ (number of models per ecosystem), $N$ (fixed data size per ecosystem, i.e., number of tokens), $n$ (data size per model), $T$ (number of iterations $t$, i.e., 10), and $\mu$ (model perplexity).
@@
In the first iteration step, the training data is identical to the generation dataset for a given model. Thereafter, it evolves in content because it is sourced from the outputs of the previous iteration: At each iteration, the $M$ generated datasets are concatenated, shuffled, and uniformly redistributed to the $M$ models for the next round. Throughout the ten iterations, the training and generation datasets for each model within an ecosystem are equal in size, $n=N/M$. For example, for two models ($D;M=2$), the sizes of the training and generation datasets are each half of the original size, $n=N/2$ (see Figure 2). In the case of the single-model ecosystem ($D;M=1$), our approach parallels the methods used in prior work on single-model collapse (e.g., Shumailov et al., 2024).
Translated to the real world, the segmentation approach broadly mirrors domain- or community-specific models (e.g., Feng et al., 2024b; Tanksley et al., 2025; Costanza-Chock, 2020) trained on distinct segments of internet data. While tailored to different communities that might use different inputs, the generated outputs are collectively stored in a single data pool, from which future training samples are drawn.
@@ 3.1. Ecosystem diversity D and model performance
Drawing from ecology, we use the HSD to measure ecosystem diversity, that is, diversity across models. The HSD quantifies the effective number of distinct, equally frequent models in a ecosystem (Roswell et al., 2021; Jost, 2006; Hill, 1973). Let $M$ denote the number of models, then we calculate HSD as:
@@
Because of equal weight across models ($w_{m}=1/M$), the diversity simplifies to the number of models in a given ecosystem (i.e. richness):
@@
Consistent with prior work (Shumailov et al., 2024), we compute each model’s perplexity on the original Wikitext2 test split to assess model performance. Perplexity reflects a model’s ability to complete the test set text sequences (lower is better) and indicates the degree of alignment between the model’s probability distribution $p_{\theta}$ and the true distribution $p_{data}$.
@@
(3) $\mu_{t}=\frac{1}{M}\sum_{m=1}^{M}\mu_{m,t}\text{ , }$
whereas $\mu_{m,t}$ corresponds to the mean perplexity across the text sequences of the original Wikitext2 test split.
To compare ecosystems across iterations, we calculate an aggregated performance. Let $\mu_{t}$ an ecosystem’s mean perplexity for a given iteration step $t$, then, the aggregated mean perplexity $\mu_{T}$ across $T$ iterations equals:
@@
Finally, for a fixed number of iterations $T$ and a set of ecosystems with varying diversity levels, we define the optimal diversity level as the one that achieves the highest aggregated performance:
(5) $D_{\mathrm{opt}}(T)=\arg\min_{D}\mu_{T}(D).$
@@ 3.2. Generation and training details
We use OPT (Zhang et al., 2022) and GPT2 (Radford et al., 2019), two open-accessible, small, and well-studied language model families used in prior work on collapse (e.g., Shumailov et al., 2024; Herel and Mikolov, 2024). For the primary approach, we use the smallest versions, OPT-125m and GPT2, with 125M and 124M parameters, respectively.
Analogous to prior work (Herel and Mikolov, 2024; Shumailov et al., 2024), we use Wikitext2 (Merity et al., 2016) as the original dataset. To reduce computational effort during training and generation, the primary approach uses a randomly sampled subset of two-fifths of the original size per ecosystem (N=0.8N=0.8M tokens), while keeping the original Wikitext2 validation and test splits for evaluation. Using a block size of 128 tokens, the sizes of the training–generation, validation, and test splits correspond to 7,500, 1,900, and 2,200 data points, respectively.
@@
To study the effect of increasing parameter size, we use OPT-350M with 350M parameters and the full Wikitext2 dataset (2.1M tokens); see Section 3.3. The analogous experiments with GPT2-Medium (355M parameters) are reported in Appendix F. All experiments are carried out using the HuggingFace Transformers library (Wolf et al., 2019) and the Python Lightning framework (Falcon and The PyTorch Lightning team, 2019).
3.3. Variations of primary setting
To understand the effect of diversity in more realistic settings, we examine how the location of the optimum shifts under two types of experimental variations: V1 examines the effect of increasing model and data sizes, and V2 examines the effect of data quality improvements through real-data incorporation and temperature sampling. We conduct the experimental variations using the OPT model family and report the scaling experiments with GPT2 in Appendix F.
V1: System scale –– model size and data size
In reality, both model and data sizes are far larger. OPT-125m contains 125M parameters, whereas current large language models (e.g., OpenAI’s ChatGPT, Google’s Gemini, Anthropic’s Claude) likely contain in the order of trillions of parameters. A similar discrepancy applies to dataset size: while Wikitext2 contains roughly 2M tokens, the datasets used for prevailing models are likely $10^{6}$ times larger. Established scaling laws suggest a power-law relationship between performance and parameter size (Kaplan et al., 2020; Hoffmann et al., 2022). However, recent work indicates that these relationships may break down under self-training conditions and that larger models and datasets may even amplify collapse (Dohmatob et al., 2024a, b). To examine ecosystem diversity through the lens of scaling laws we introduce three variations of the primary setting: (1) a larger model (OPT-350m, with 350M parameters), (2) the full Wikitext2 training set (2.1M tokens), and (3) the combination of both larger model and dataset sizes. We conduct these experiments for M=M=1, 2, 4, 8, and 16 models.
V2: Training data quality –– real-data incorporation and sampling temperature
In realistic self-training loops, artificial data does not replace but rather accumulates alongside real (non-model-generated) data, improving the quality of the training set and mitigating collapse (Shumailov et al., 2024; Gerstgrasser et al., 2024; Xing et al., 2025). Similarly, prevailing language models are not deterministic but incorporate randomness that diversifies outputs and can mitigate model collapse (Wright et al., 2025; Schaffelder and Gatt, 2025; Zhu et al., 2025). A common mechanism is temperature sampling, which increases output diversity by flattening token-sampling probabilities (Chung et al., 2023; Alihosseini et al., 2019). To test these factors, we introduce an experimental variation that replaces 10% of the model-generated data with randomly sampled original Wikitext2 training data at each iteration (following Shumailov et al. (2024)), as well as three additional variations using sampling temperatures of $\tau=0.5$, $1.0$, and $2.0$. As a reminder, the primary experiment uses deterministic generation with $\tau=0$. When temperature sampling is applied, all remaining sampling parameters follow the Transformers default settings, corresponding to top_k = 50 and top_p = 1.0.
@@ 4. Results
We conducted a series of experiments to examine how diversity across models affects performance when models are trained on their collective outputs. Under a fixed data size and compute budget, we find that the optimal degree of diversity increases monotonically with the number of self-training iterations:
(6) $T_{1}<T_{2}\Longrightarrow D_{opt}(T_{1})\leq D_{opt}(T_{2}),$
where the optimal level is defined as the ecosystem achieving the highest aggregated performance (i.e., lowest aggregated mean perplexity). In other words, as the number of self-training iterations increases, maximizing overall performance requires segmenting the data across more models. Correspondingly, the benefits of diversity relative to the most homogeneous setting grow with the number of iterations.
In the primary approach, the single-model setting ($D=1$) performs best only in the short term (considering fewer than four iterations for OPT-125M and fewer than three iterations for GPT2, see Figure 3) Figure 4 shows the corresponding mean perplexity trajectories for ecosystems consisting of 1, 2, 4, and 16 models trained on equal-sized, non-overlapping segments of model-generated data across 10 iterations. Ecosystems with lower diversity ($D;M=1,2$) perform well initially but exhibit rapidly increasing perplexity as the number of iterations grows. In contrast, the most diverse ecosystem ($M=16$) starts with relatively high perplexity but improves over time. Across the full 10 iterations, a diversity level of four ($D;M=4$) yields the lowest aggregated mean perplexity for both model families, with a stronger effect for OPT-125M. The least diverse ecosystem ($D=1$) performs best only at $t=0$, when ecosystems differ solely in sample size per model ($n=N/M$), as all models are trained on the original data [^5] The perplexity progression, including the perplexity of the original models, is shown in Figure 8 in the appendix..
Figure 10 in the Appendix extends the analysis to 20 iterations for OPT-125M, showing that the optimal diversity level continues to increase beyond 10 iterations.
Figure 3. The optimal level of ecosystem diversity (Hill–Shannon diversity $D$) increases monotonically with the number of self-training iterations $T$. The most homogeneous setting ($D=1$, green) performs best only in the short term ($T\leq 3$ for OPT-125M and $T\leq 2 f$or GPT2). Thus, diversity $D$ mitigates model collapse. As $T$ grows, the data must be segmented across models to reach the diversity level that maximizes overall performance on the original test set ($D_{opt}(T)={\mathrm{argmin}_{D}}\,\mu_{T}(D)$). The corresponding performance trajectories are presented in Figure 4. Colored symbols correspond to the diversity levels shown in Figure 2 1, 2, 4, and 16 (e.g., orange denotes $D=2$). Intermediate levels are omitted, as the key result is the monotonic increase in diversity benefits with $T$. An extended range of up to 20 iterations is shown in Figure 10 in the Appendix.
@@
(c) Aggregated perplexity after 10 iterations
Figure 4. Ecosystem diversity mitigates collapse. (a, b) Comparison of ecosystems’ mean perplexity with increasing diversity ($D=M$) at fixed dataset size across ten iteration steps. While low-diversity ecosystems ($D=1,2$) exhibit strong performance decay, the high-diversity ecosystem ($D=16$) starts with relatively high perplexity but improves over $t$ . The shaded region is a combined standard deviation across the models and data points. (c) Aggregated perplexity. Both model families exhibit an optimal level of diversity of 4 for overall performance across 10 iterations. Colors correspond to the diversity levels shown in Figure 2 (e.g., orange denotes $M;D=2$).
(a) Optimal diversity levels.
(b) Aggregated perplexities after 10 iterations.
(c) Mean perplexities for M;D=1
(d) Mean perplexities for M;D=16
[…]
(a) Optimal diversity levels.
(b) M;D=1
(c) M;D=4
[…]
(a) Optimal diversity levels.
(b) M;D=1
(c) M;D=16
[…]
4.1. V1: Model size and data size
The benefits of diversity increase as both model size (from 125M to 350M parameters) and dataset size (from 0.8M to 2.1M tokens) scale (Figure 5). Holding dataset size constant, larger models exhibit a steeper increase in the optimal diversity level as the number of iterations grows. A similar pattern appears when increasing dataset size for the larger model, whereas for the smaller model dataset size has only a minor effect on the optimal diversity level. Below, we highlight several notable performance effects relative to the primary setting (125M-parameter model and 0.8M tokens dataset).
At iteration step $t=0$, when models are trained on the original dataset, increasing either model or dataset size improves performance in line with established scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). The most pronounced effects of larger models is the amplified collapse of low-diversity ecosystems ($D\leq 2$), leading to the observed shift in the optimal diversity level. While the effects of dataset size are less pronounced, two patterns emerge: increasing dataset size for the the larger model improves performance primarily in the most homogeneous ecosystem ($D=1$), whereas for the smaller model it improves performance primarily in the most diverse setting ($D=16$).
4.2. V2: Real-data incorporation and sampling temperature
Improving data quality by incorporating 10% real data (Figure 6) or by applying temperature sampling with $\tau=1.0$ and $\tau=2.0$ (Figure 7) generally improves model performance. Because both interventions primarily benefit the most homogeneous ecosystem ($D=1$), performance differences across diversity levels shrink relative to the primary setting. However, since the positive effect of ecosystem diversity exceeds these improvements, both interventions produce a similar trajectory of the optimal diversity level over iterations as observed in the primary setting. We highlight several notable performance effects relative to the primary setting below.
The impact of incorporating real data at each iteration increases as diversity decreases, reflecting the larger per-model dataset size ($n=N/M$). Temperature sampling can also affect data quality by diversifying generated outputs. For temperatures of $\tau=1.0$ and $\tau=2.0$ in the most homogeneous ecosystem ($D=1$), perplexity increases during early iterations but declines by a similar magnitude later in the loop, resulting in only marginal overall improvement. In more diverse ecosystems, $\tau=1.0$ produces slightly negative effects, while $\tau=2.0$ leads to strongly negative effects. By contrast, a temperature of $\tau=0.5$ shows almost no effect relative to the primary setting for any fixed diversity level.
@@ 5. Theoretical interpretation
Across experiments, our study of ecosystem evolution reveals a monotonic increase of the optimal diversity level with number of self-training iterations $T$. For a fixed number of 10 iterations, we find a U-shaped relationship between diversity and collapse. Considering that dataset size per model $n=N/M$ decreases as diversity increases, the observed increase in optimum diversity level indicates a complex interplay between opposing effects of diversity and scale with $T$. The complexity grows when considering the results obtained with increased scale (Section 4.1): When model size increases, we observe stronger collapse in low-diversity ecosystems ($D\leq 2$) and weaker collapse in high-diversity ecosystems ($D\geq 4$). Increasing dataset size generally improves performance for OPT-125m but worsens performance for OPT-350m, except in the $D=1$ ecosystem. In this section we propose a theoretical explanation of the underlying effects.
Dohmatob et al. (2024a) theorize that increasing model and/or dataset size amplifies model collapse once data quality becomes sufficiently low. Building on this theory, we can assume that low-diversity ecosystems ($D\leq 2$) tend to operate in a low data-quality regime, leading to stronger collapse as model size increases, while the opposite holds for high-diversity ecosystems ($D\geq 4$). This raises the question of how diversity helps preserve high data quality, given that diverse ecosystems start with lower model performance at iteration step $t=0$. To address this, we introduce effective data quality (EDQ), which builds on the framework of Dohmatob et al. (2024a). We then use this hypothesis to explain our empirical results in detail.
Effective data quality EDQ
Training data quality reflects the discrepancy between $\hat{p}{data}$ and pdatap{data} (Dohmatob et al., 2024a; Wang et al., 2025b), for example as measured by (Shumailov et al., 2024; Bertrand et al., 2024): $\delta(\hat{p}_{data})=\mathbb{W}_{2}(\hat{p}_{data},p_{data})$, whereas $\mathbb{W}_{2}$ refers to the Wasserstein-2 distance between the two distributions [^6] The probability distributions are illustrated in Figure 11 in the appendix.. For an original (non–model-generated) dataset, all examples lie within the support of the true distribution, and this discrepancy approaches zero as dataset size approaches infinity. In contrast, due to models’ internal errors (functional expressivity and functional approximation errors (Shumailov et al., 2024)), resampled artificial datasets can contain less information about the true distribution and can even include examples outside its support. Moreover, depending on the data received in previous training steps, not every in-distribution example provides new or useful information to a model (Feng et al., 2024a; Sorscher et al., 2023) and further training on a dataset similar to the initial training set risks performance decay via overfitting (Tirumala et al., 2022). Effective fine-tuning methods suggest that optimal training data is not merely data that reflects the true distribution, but data that most efficiently moves model weights toward that distribution (Ouyang et al., 2022; Liu et al., 2025; Ji et al., 2024; Deb et al., 2025; Antonello et al., 2021). Hence, optimal data must account for the model’s existing distribution, which approximates a mixture of previous training distributions, i.e., ${p}_{\theta(t)}\approx\sum_{i=0}^{t-1}\hat{p}_{data(i)}$. We define EDQ at iteration $t$ as the difference between the training–truth discrepancy and the model–truth discrepancy [^7] Because we introduce EDQ primarily to explain our empirical results, we rely on a simple scalar metric to capture distributional discrepancy and ignore additional factors required for optimal model weights updates, such as the direction of the distributional shift.:
(7) $\delta_{EDQ}(\hat{p}_{data},{p}_{\theta)}:=\mathbb{W}_{2}(\hat{p}_{data(t)},p_{data})-\mathbb{W}_{2}({p}_{\theta(t)},p_{data}),$
where lower values of $\delta_{EDQ}$ indicate higher EDQ. EDQ reflects the model-dependent informativeness of the given training data. Using $\delta_{EDQ}$, we can now state our main hypothesis under finite training data and compute:
Hypothesis: EDQ, informal
  1. (1)
    High-EDQ regime: If $\mathbb{W}_{2}(\hat{p}_{data(t)},p_{data}))$ sufficiently lower than $\mathbb{W}_{2}({p}_{\theta(t)},p_{data})$, additional training improves performance.
  2. (2)
    Low-EDQ regime : If $\mathbb{W}_{2}(\hat{p}_{data(t)},p_{data}))$ sufficiently greater than $\mathbb{W}_{2}({p}_{\theta(t)},p_{data})$, additional training degrades performance.
  3. (3)
    Medium-EDQ regime: If $\mathbb{W}_{2}(\hat{p}_{data(t)},p_{data}))\approx\mathbb{W}_{2}({p}_{\theta(t)},p_{data})$, additional training neither significantly improves nor degrades model performance.
This hypothesis has two implications for recursively trained models. First, performance gains or losses depend on EDQ, regardless of whether the data is generated by humans or models (Wang et al., 2025b; Feng et al., 2024a), or on its diversity per se. Nevertheless, both the proportion of artificial data and data diversity can shape the EDQ. Second, when operating in a low-EDQ regime, neural scaling laws invert: performance decay increases with system scale, consistent with prior findings (Dohmatob et al., 2024a; Herel and Mikolov, 2024; Wright et al., 2025). Neural Scaling laws state that test loss depends on model size and dataset size, following a power-law relationship (Kaplan et al., 2020; Hoffmann et al., 2022).
Diversity and EDQ
The EDQ hypothesis helps explain the impact of ecosystem diversity on model performance. First, in our segmentation approach, sample size per model decreases as diversity increases; thus, once models operate in a low-EDQ regime, greater diversity becomes beneficial. This relationship is consistent with prior work (Dohmatob et al., 2024a) and reflects a trade-off in long-term considerations of AI ecosystems between large-scale under high-EDQ and small-scale under low-EDQ. Second, the distribution of an artificial dataset closely resembles the distribution of the model from which it is sampled. If a model had zero sampling error, these distributions would be identical. However, under our hypothesis, data whose distribution closely mirrors the model itself has low EDQ. In contrast, in a diverse ecosystem, different models represent different components of the true distribution. As a result, the EDQ of redistributed outputs can be substantially higher than in a setting where models only see their own output: data that is not useful for the model that generated it may still be useful for another model within the same ecosystem (see Illustration 12 in the Appendix). This introduces a second trade-off for maximizing EDQ: between larger datasets per model when the data are fresh to the model ($t=0$) and higher EDQ through diversity in later iterations.
We now use the proposed theory to interpret performance dynamics across different stages of self-training iterations.
Iteration step $t=0$ At $t=0$, all models are trained on subsets of real data and therefore operate in a high-EDQ regime. EDQ increases with sample size per model $n=N/M$, which decreases as the number of models $M$ increases. Because default scaling laws apply in this regime, performance decreases with diversity and improves with system scale (see Section 4.1).
Later iterations In later iterations, more diverse ecosystems can exhibit lower performance loss (at $t=1$) and/or greater performance gains because (i) the exchange of artificial data among models increases EDQ, and (ii) once models enter the low-EDQ regime, larger per-model sample sizes $n$ become harmful. This explains the sharp decay observed for $D=1$ in Figure 4, which is further amplified by larger models (Figure 5). Performance trajectories after $t=0$ illustrate that models in diverse ecosystems receive data of higher EDQ than models in homogeneous ecosystems, ultimately yielding superior long-term performance, despite lower model performance at the beginning.
Post hoc sensitivity analyses of individual factors that may influence the performance dynamics observed in the evolution of AI ecosystems support our EDQ hypothesis (see Appendix H). The results indicate that neither (model-independent) data quality, training data size, the number of models per ecosystem, nor overfitting can fully account for the observed performance decay or the differences across ecosystems.